Use the midpoint approximation method to find the area under the curve 𝑓(𝑥)=1−𝑥2 over

the interval [−1,1].


1
Expert's answer
2021-09-27T18:18:25-0400

The midpoint rule (also known as the midpoint approximation) uses the midpoints of a subinterval:

  

"\begin{aligned}

&\int_{a}^{b} f(x) d x \\

&\approx \Delta x\left(f\left(\frac{x_{0}+x_{1}}{2}\right)+f\left(\frac{x_{1}+x_{2}}{2}\right)+f\left(\frac{x_{2}+x_{3}}{2}\right)+\cdots+f\left(\frac{x_{n-2}+x_{n-1}}{2}\right)\right. \\

&\left.+f\left(\frac{x_{n-1}+x_{n}}{2}\right)\right) \\

\end{aligned}"

 

where  "\Delta x=\frac{b-a}{n}"

We have that "f(x)=1-x^{2}, a=-1, b=1 ," and "n=4"

Therefore,  "\Delta x=\frac{1-(-1)}{4}=\frac{1}{2}"

Divide the interval "[-1,1]"  into  "n=4"  subintervals of the length  "\Delta x=\frac{1}{2}"  with the following endpoints:  "a=-1,-\frac{1}{2}, 0, \frac{1}{2}, 1=b"

Now, just evaluate the function at the midpoints of the subintervals.

  

"\begin{aligned}

&f\left(\frac{x_{0}+x_{1}}{2}\right)=f\left(\frac{-1-\frac{1}{2}}{2}\right)=f\left(-\frac{3}{4}\right)=\frac{7}{16}=0.4375 \\

&f\left(\frac{x_{1}+x_{2}}{2}\right)=f\left(\frac{-\frac{1}{2}+0}{2}\right)=f\left(-\frac{1}{4}\right)=\frac{15}{16}=0.9375 \\

&f\left(\frac{x_{2}+x_{3}}{2}\right)=f\left(\frac{0+\frac{1}{2}}{2}\right)=f\left(\frac{1}{4}\right)=\frac{15}{16}=0.9375 \\

&f\left(\frac{x_{3}+x_{4}}{2}\right)=f\left(\frac{\frac{1}{2}+1}{2}\right)=f\left(\frac{3}{4}\right)=\frac{7}{16}=0.4375

\end{aligned}"

Finally, just sum up the above values and multiply by  "\Delta x=\frac{1}{2} :"

 

"\frac{1}{2}(0.4375+0.9375+0.9375+0.4375)=1.375"